OpenStudy (wcrmelissa2001):

Help please thanks :D Linear inequalities: When do we have to consider cases (e.g. when x<1) and when do we not have to? Coz I was just doing this question which apparently needs cases but I don't understand why. 2/(3(x+1)) + 1/(3(x+2))≤1/(x+1)

OpenStudy (wcrmelissa2001):

and also how do you solve the above inequality i thought i could but i can't apparently

OpenStudy (phi):

the technique is to compare the expression to zero it's always "safe" to add or subtract a quantity (but you can get into trouble if you multiply or divide). So write the problem as \[ \frac{2}{3(x+1)} + \frac{1}{3(x+2)} - \frac{1}{x+1} \le 0 \]

OpenStudy (wcrmelissa2001):

yup then find the common denominator?

OpenStudy (phi):

yes, combine the fractions

OpenStudy (wcrmelissa2001):

x+3/(x+1)(x+2)(x+1)

OpenStudy (phi):

I get something else

OpenStudy (wcrmelissa2001):

which is?

OpenStudy (phi):

what are you using for a common denominator ?

OpenStudy (wcrmelissa2001):

initially after combining it's something like (3x-9)/3(x+1)(x+2)(x-1) = x-3/(x+1)(x+2)(x-1)

OpenStudy (phi):

I would use 3(x+1)(x+2) as the common denominator

OpenStudy (wcrmelissa2001):

ok

OpenStudy (wcrmelissa2001):

any particular reason?

OpenStudy (phi):

if you do, multiply the first fraction by (x+2)/(x+2)

OpenStudy (wcrmelissa2001):

if 3(x+1)(x+2) as the common denominator what's the numerator?

OpenStudy (phi):

the reason is that is the simplest common denominator. ***what's the numerator? *** multiply the first fraction by (x+2)/(x+2)

OpenStudy (wcrmelissa2001):

as in x+3 or 3x-9?

OpenStudy (phi):

I would never distribute in these problems. we want to keep things factored the first fraction becomes \[ \frac{2}{3(x+1)} \cdot \frac{(x+2)}{(x+2)} \]

OpenStudy (wcrmelissa2001):

isn't it 2/3(x-1) not x+1?

OpenStudy (phi):

maybe you should post a screen shot? I am looking at what you posted in the question up top. If there are typo's , this will be very confusing...

OpenStudy (wcrmelissa2001):

erm it's hardcopy but i'll type it out \[\frac{ 2 }{ 3(x-1) }+\frac{ 1 }{ 3(x+2) }≤\frac{ 1 }{ x+1 }\]

OpenStudy (phi):

oh, that is different from your question where it's (x+1) in the first fraction

OpenStudy (wcrmelissa2001):

OH CRAP SORRY

OpenStudy (phi):

so your answer up above is mostly correct (except for x+1 twice and no x-1) ***x+3/(x+1)(x+2)(x+1)*** you mean \[ \frac{(x+3)}{(x-1)(x+1)(x+2) }\]

OpenStudy (phi):

so the problem is \[ \frac{(x+3)}{(x-1)(x+1)(x+2) } \le 0 \]

OpenStudy (wcrmelissa2001):

ok yup :D

OpenStudy (phi):

we need that fraction to be 0 or negative

OpenStudy (phi):

it's zero for x= -3. make a note of that

OpenStudy (wcrmelissa2001):

that's where cases come in yay. looks like there'll be a lot?

OpenStudy (phi):

yes, it looks a bit messy. I would plot a number line, and figure out the signs of each factor for each region of the number line.

OpenStudy (wcrmelissa2001):

ok so am i right to say that er the cases are something like ok wait. gimme a moment sorry

OpenStudy (wcrmelissa2001):

case 1: x+3≤0 and everything else ≥0 x>1 case 2: x-1≤0 and everything else more than 0 -1<x<1 case 3: x+2≤0 everhting else ≥0 -3<x<2 or x>1 case 4: x+1≤0 -3<x<-1 or x>1??

OpenStudy (wcrmelissa2001):

sorry forgot the equal sign for inequality

OpenStudy (phi):

notice we can't let the bottom factors be 0 (no divide by 0)

OpenStudy (wcrmelissa2001):

ahh

OpenStudy (wcrmelissa2001):

case 4 should be x>-1 i think

OpenStudy (phi):

graphically